A painting in an art gallery has height 3 m and is hung so that its lower edge is about 1 m above the eye of an observer. How far from the wall should the observer stand to get the best view? (i.e. the observer wants to maximize the angel θ subtended at the eye by the painting


1
Expert's answer
2020-11-17T06:17:57-0500

Let man is at a distance of "x" from the wall.

Let angle made made at lower and upper end be "\alpha" and "\theta" respectively.

for sake of simplicity, Let height of the painting is "h=3m"

and height of the lower end from the ground is, "d=1m"


Then,

"tan(\theta + \alpha) = \frac{h+d}{x}"

Differentiating both sides,

"sec^2(\theta + \alpha)(\theta'(x) + \alpha'(x))=-\frac{h+d}{x^2}"


"\theta'(x) + \alpha'(x)=-\frac{(h+d)cos^2(\theta + \alpha)}{x^2}"


Putting value of "cos(\theta+\alpha) = \frac{x}{\sqrt{x^2+(d+h)^2}}"


then,

"\theta'(x) + \alpha'(x)=-\frac{(h+d)x^2}{x^2(x^2+(d+h)^2)} = -\frac{h+d}{x^2+(d+h)^2}" (1)


Now,

"tan \alpha = \frac{d}{x} \implies \alpha =tan^{-1} \frac{d}{x}"


Differentiating it, we get,

"\alpha'(x) = \frac{x^2}{d^2+x^2}\frac{-d}{x^2} = -\frac{d}{d^2+x^2}" (2)


From (1) and (2),

we get,

"\theta'(x) = -\frac{h+d}{x^2+(d+h)^2}- \alpha'(x)
= \frac{d}{d^2+x^2}-\frac{h+d}{x^2+(d+h)^2}"

To maximize the area, put

"\theta'(x) = -\frac{h+d}{x^2+(d+h)^2}- \alpha'(x)
= \frac{d}{d^2+x^2}-\frac{h+d}{x^2+(d+h)^2} = 0"

Solving it, we get,

"d(h+d)^2+dx^2=(h+d)(d^2+x^2)"

"x = \sqrt{d(d+h)}"


Putting value of d and h in above equation,

"x = 2 m"